# From Intention To Implementation: Automating Biomedical Research via LLMs

### v1

@@ 5 Experiment
    How does BioResearcher perform in automating the entire biomedical research process?
@@
    How does the performance of the query generator and filter agents compare to that of humans?
  3. 3.
    How effective is our hierarchical report generation method?
  4. 4.
    What are the impacts of our hierarchical learning approach, as well as the reviewer and report analyst agents, on the experimental design?
@@ 5.1 Experiment Setup
We evaluate BioResearcher by comparing it with three well-known agent systems. (1) ReAct (Yao et al., 2023), which integrates reasoning and action within LLMs to effectively manage complex reasoning and decision-making tasks. (2) Plan-and-Execute (Wang et al., 2023c), which utilizes an iterative framework to accomplish tasks through sequential planning and execution. (3) A naive RAG-based LLM system, which employs a naive RAG module to search for relevant content and generates the answer in a single step. Section A provides a detailed description of the implementation of these baseline systems.
@@ 5.2.1 Results
Table 3 presents the average scores for protocols generated by different systems, including Completeness, Level of Detail, Correctness, Logical Soundness, and Structural Soundness. We also calculate the overall performance, the average number of sentences per step ($l_{\text{steps}}$), and the average number of steps per protocol ($n_{\text{total steps}}$). (1) The results highlight BioResearcher’s exceptional performance in Completeness, Level of Detail, and Correctness, surpassing the best baseline by 0.220 (56.1%), 0.137 (17.9%), and 0.457 (93.8%), respectively. (2) Our system is comparable to the best baseline in Logical Soundness and Structural Soundness, both exceeding 0.9, consistently maintaining a high standard in these areas. (3) Furthermore, the protocols generated by BioResearcher have significantly more sentences per step and steps per protocol, 5.9×5.9$\times$5.9 × and 4.8×4.8$\times$4.8 × greater than the average performance of different baselines, respectively, indicating the generation of more detailed and comprehensive protocols. (4) The comparative analysis of the three baselines indicates that the Naive RAG LLM outperforms the other two iterative agent systems. This superiority can be attributed to the fact that, in the long-context environment characterized by iteration and lengthy retrieved information, the summarization and integration capabilities of the latter two systems degrade, resulting in the loss of substantial amounts of valuable information in the final generated protocol. In contrast, we design tailored workflows to effectively integrate results from multiple modules and steps, thereby preserving critical information throughout the integration process. A case comparison of protocols from the four systems is illustrated in Figure 10.
@@ 5.2.2 The Quality of LLM Evaluation
Table 4: Performance of LLM and Human Experts’ Evaluation
@@ 5.4 Performance of Report Generation
To test the impact of our hierarchical report generation method, we compare its reports against those produced by ReAct, Plan-and-Execute, and a single-step LLM approach. Specifically, we standardized 20 papers into an experimental report format using each method. None of the methods are equipped with any additional tools. We prompt an LLM to evaluate the reports across four dimensions: logical soundness, level of detail, consistency with the original paper, and readability. The model assigns a score ranging from 1 to 5 for each dimension, with the scoring criteria detailed in Table 6. The final score is calculated as the average across these four dimensions. Figure 9 presents a box plot of the scores for reports generated by the four different methods.
@@ Appendix A Details of Baselines Implementations
In Section 5.2, due to the NCBI databases’ limitations in semantic similarity searches, directly using the user’s input objective poses challenges in retrieving relevant papers and datasets. Thus, we provide the naive RAG-based LLM system with unfiltered papers and datasets from our Search module, enabling it to extract relevant datasets and content to generate a comprehensive protocol and corresponding code in a single step.
@@
We implement all baselines using the AutoGen framework (Wu et al., 2023) and adopt OpenAI’s text-embedding-ada-002 as the embedding model in RAG applications.
@@ Appendix B Details of Research Objectives
Table 10: 10 Research Objectives Used in The Evaluation of Search module.
@@
8 To reveal the transcriptional patterns of HER2 through a pan-cancer analysis of HER2 indices, facilitating more precise selection of patients suitable for HER2-targeted therapy.
@@
10 To reveal the heterogeneity of T cell exhaustion (TEX) in the tumor microenvironment (TME) of different cancer types, depict the hierarchical functional dysfunction process of T cell exhaustion through pan-cancer analysis, and investigate its association with prognosis and immune therapy efficacy.
Table 11: 15 Research Objectives Used in The Evaluation of Literature Processing, Experimental Design Module, and LLM Reviewers.
@@ Appendix C Cost Analysis
Table 13: Time comparison between Human and BioResearcher with GPT-4o
@@
Figure 10 shows a comparison of a part of the experimental protocols generated by four different systems: BioResearcher, Plan-and-execute, React, and Naïve RAG-based LLM. Notably, the protocol produced by BioResearcher is both detailed and accurate. In contrast, the protocols generated by the other systems lack essential details, rendering them impractical for implementation.
Figure 10: A comparison of BioResearcher, Plan-and-Execute, React, and Naïve Rag in generating protocols.

### v2

@@ 5 Experiment
    RQ1: How does BioResearcher perform in automating the entire biomedical research process?
@@
    RQ2: How does each component of BioResearcher perform in their specific sub-tasks, including the search module (RQ2.1), report generation module (RQ2.2), and experimental design module (RQ2.3)?
@@ 5.1 Experiment Setup
We evaluate BioResearcher by comparing it with three well-known agent systems. (1) ReAct (Yao et al., 2023), which integrates reasoning and action within LLMs to effectively manage complex reasoning and decision-making tasks. (2) Plan-and-Execute (Wang et al., 2023c), which utilizes an iterative framework to accomplish tasks through sequential planning and execution. (3) RAG, which employs a naive Retrieval-Augmented Generation (RAG) module to search for relevant content and generates the answer in a single step. Section A provides a detailed description of the implementation of these baseline systems.
@@ 5.2.1 Results
Table 3 presents the average scores for protocols generated by different systems, including Completeness, Level of Detail, Correctness, Logical Soundness, and Structural Soundness. We also calculate the overall performance, the average number of sentences per step ($l_{\text{steps}}$), and the average number of steps per protocol ($n_{\text{total steps}}$). (1) The results highlight BioResearcher’s exceptional performance in Completeness, Level of Detail, and Correctness, surpassing the best baseline by 0.220 (56.1%), 0.137 (17.9%), and 0.457 (93.8%), respectively. (2) Our system is comparable to the best baseline in Logical Soundness and Structural Soundness, exceeding 0.9, consistently maintaining a high performance in these aspects. (3) Furthermore, the protocols generated by BioResearcher have significantly more sentences per step and steps per protocol, 5.9×5.9$\times$5.9 × and 4.8×4.8$\times$4.8 × greater than the average performance of different baselines, respectively, indicating the generation of more detailed and comprehensive protocols. (4) The comparative analysis of the three baselines indicates that RAG outperforms the other two iterative agent systems. This superiority can be attributed to the fact that, in the long-context environment characterized by iteration and lengthy retrieved information, the summarization and integration capabilities of the latter two systems degrade, resulting in the loss of substantial amounts of valuable information in the final generated protocol. In contrast, we design tailored workflows to effectively integrate results from multiple modules and steps, thereby preserving critical information throughout the integration process. A case comparison of protocols from the four systems is illustrated in Figure 10.
@@ 5.2.2 The Quality of LLM Evaluation
Table 4: Performance of LLM and human experts’ evaluation.
@@ 5.4 Performance of Report Generation
To test the impact of our hierarchical report generation method, we compare its reports against those produced by ReAct, Plan-and-Execute, and a naive single-step LLM approach. Specifically, we standardized 20 papers into an experimental report format using each method. None of the methods are equipped with any additional tools. We prompt an LLM to evaluate the reports across four dimensions: logical soundness, level of detail, consistency with the original paper, and readability. The model assigns a score ranging from 1 to 5 for each dimension, with the scoring criteria detailed in Table 6. The final score is calculated as the average across these four dimensions. Figure 9 presents a box plot of the scores for reports generated by the four different methods.
@@ Appendix A Details of Baselines Implementations
In Section 5.2, due to the NCBI databases’ limitations in semantic similarity searches, directly using the user’s input objective poses challenges in retrieving relevant papers and datasets. Thus, we provide the RAG system with unfiltered papers and datasets from our Search module, enabling it to extract relevant datasets and content to generate a comprehensive protocol and corresponding code in a single step.
@@
We implement all baselines using the AutoGen framework Wu et al. (2023) and adopt OpenAI’s text-embedding-ada-002 as the embedding model in RAG applications.
@@ Appendix B Details of Research Objectives
In this section, we outline the specific research objectives utilized in our experiments. Table 9 presents eight objectives employed for end-to-end automation experiments. These objectives are previously unmet and are gathered from a biomedical research laboratory. Table 10 displays ten objectives sourced from publicly available papers for evaluating the search module. Table 11 presents fifteen objectives for assessing experimental design; these are also collected from published papers, but their publication dates are later than the cutoff for GPT-4o training data.
@@
Table 10: 10 research objectives used in the evaluation of the Search module.
@@
8 To reveal the transcriptional patterns of HER2 through a pan-cancer analysis of HER2 indices, facilitating a more precise selection of patients suitable for HER2-targeted therapy.
@@
10 To reveal the heterogeneity of T cell exhaustion (TEX) in the tumor microenvironment (TME) of different cancer types, depict the hierarchical functional dysfunction process of T cell exhaustion through pan-cancer analysis and investigate its association with prognosis and immune therapy efficacy.
Table 11: 15 research objectives used in the evaluation Experimental Design.
@@ Appendix C Cost Analysis
Table 13: Time comparison between human and BioResearcher
@@
Figure 10 shows a comparison of a part of the experimental protocols generated by four different systems: BioResearcher, Plan-and-execute, React, and RAG. Notably, the protocol produced by BioResearcher is both detailed and accurate. In contrast, the protocols generated by the other systems lack essential details, rendering them impractical for implementation.
Figure 10: A comparison of BioResearcher, Plan-and-Execute, React, and RAG in generating protocols.